OpenStudy (anonymous):

You are going to purchase a new car, but being a responsible consumer means doing a little bit of research first. First, you find the vehicle you are purchasing and its price. Vehicle: Chevy Volt Price: $39,145 Current interest rate: 3%

OpenStudy (anonymous):

Using the function A(t)=P(1+r/n)^nt, create the function that represents your new car loan that is compounded monthly. The principle will be the price of the vehicle you selected, not how much you are putting down.

OpenStudy (phi):

monthly means 12 times a year. n is 12 r is the year rate as a decimal. change 3% to a decimal t will the the years of the loan. 3,4 or even 5 years

OpenStudy (anonymous):

So A(t) = 39,145(1+ 0.3/12)^(12t) ?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

0.3 is 3/10 or 30/100 or 30% don't you want 3% ?

OpenStudy (anonymous):

Yeah A(t) = 39,145(1+ 0.03/12)^(12t)

OpenStudy (anonymous):

Now what do we do?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

The problem said to create a function that represents the loan. It looks like you did so, though I'll observe that the function represents the loan if no payments are made, and you probably won't reach the end of the loan period with the car still in your possession if you go that route :-)

OpenStudy (anonymous):

Thanks, can you help me with the last two parts of this?

OpenStudy (whpalmer4):

What are the last two parts of this?